Find two consecutive odd integers such that their product is 71 more than 7 times their sum
Answer by jorel555(1290)

let n be your first integer; your second integer is n+2:
n(n+2)=7(n+n+2)+71
n^2+2n=14n+85
n^2-12n-85=0
(n-17)(n+5)=0
n=17
n+2=19!!!!!!!!